### v2

# On the kissing number of the cross-polytope

Abstract
A new upper bound $\kappa_{T}(K_{n})\leq 2.9162^{(1+o(1))n}$ for the translative kissing number of the $n$-dimensional cross-polytope $K_{n}$ is proved, improving on Hadwiger’s bound $\kappa_{T}(K_{n})\leq 3^{n}-1$ from 1957. Furthermore, it is shown that there exist kissing configurations satisfying $\kappa_{T}(K_{n})\geq 1.1637^{(1-o(1))n}$, which improves on the previous best lower bound $\kappa_{T}(K_{n})\geq 1.1348^{(1-o(1))n}$ by Talata. It is also shown that the lattice kissing number satisfies $\kappa_{L}(K_{n})<12(2^{n}-1)$ for all $n\geq 1$, and that the lattice $D_{4}^{+}$ is the unique lattice, up to signed permutations of coordinates, attaining the maximum lattice kissing number $\kappa_{L}(K_{4})=40$ in four dimensions.
1 Introduction
The translative kissing number problem asks for the maximum number of non-overlapping translates of a given convex body that can touch a central one. The kissing number of the $n$-dimensional $\ell^{2}$-ball $B_{n}=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}^{2}\leq 1\}$ is known in dimensions 11114444, $8$ and $24$ [1, 2, 3, 4], and the asymptotic bounds $2^{0.2075{(1+o(1))n}}\leq\kappa_{T}(B_{n})\leq 2^{0.401(1+o(1))n}$ follow from the works of Wyner [5], Kabatiansky and Levenshtein [6] and Levenshtein [3]. The kissing number problem for convex bodies different from $B_{n}$ are equally interesting and important. See [7] for a survey on kissing numbers of convex bodies and related problems.
In this article, we focus on the kissing number problem for the $n$-dimensional cross-polytope $K_{n}=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}|x_{i}|\leq 1\}$, i.e., the convex hull of the $n$$\pm$ pairs of standard basis vectors in $\mathbb{R}^{n}$. The cross-polytope is, together with the cube and the simplex, the only regular polytope in dimension $n\geq 5$. The exact value of $\kappa_{T}(K_{n})$ is known only in dimensions 11113333 where it is $2$, $8$ and $18$ [8], respectively. In each of these dimensions, the kissing configuration consists of the shortest non-zero vectors of a lattice – the $D_{n}$ lattice, defined as the set of elements of $\mathbb{Z}^{n}$ whose coordinates sum up to an even number. In dimension four, the shortest vectors of the $D_{4}$ lattice gives a kissing configuration with $32$ vectors. However, the lattice $D_{4}$ together with the translate $D_{4}+(\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2})$ produce a kissing configuration with $40$ vectors, which we show is optimal in Theorem 2. Moreover, the optimal lattice kissing configuration is unique up to equivalence (see Section 2 for definitions).
Using a volume argument, Hadwiger [9] proved that the bound $\kappa_{T}(C)\leq 3^{n}-1$ holds for any $n$-dimensional convex body, a bound which remains the best general upper bound for convex bodies. In Section 4, we use a variant of Blichfeldt’s enlargement method [10] to show that
Theorem 1.
$\kappa_{T}(K_{n})\leq 2.9162^{(1+o(1))n}$ as $n\to\infty$.
This seems to be the first improvement on Hadwiger’s bound for $\ell^{p}$-balls different from $B_{n}$ (note that $\kappa_{T}(Q_{n})=3^{n}-1$, where $Q_{n}$ is the $n$-cube). The lattice kissing number bound $\kappa_{L}(C)\leq 2(2^{n}-1)$ for strictly convex $C$ is a classical bound by Minkowski [11], see also [12, Theorem 30.2]. While $K_{n}$ is not strictly convex, it is possible to obtain an equally strong inequality for $\kappa_{L}(K_{n})$, up to a constant factor. See Section 3.
2 Preliminaries
In this section, we recall basic facts about convex bodies and kissing configurations. By a convex body $C$, we mean a compact and convex subset of $\mathbb{R}^{n}$ with non-empty interior. The set of centrally symmetric convex bodies in $\mathbb{R}^{n}$ is denoted $\mathcal{K}_{n}^{0}$. The symmetry group of $C\in\mathcal{K}_{n}^{0}$ is defined as the group $\mathcal{G}_{C}=\{T\in O_{n}(\mathbb{R}):T(C)=C\}$, where $O_{n}(\mathbb{R})$ is the orthogonal group consisting of distance-preserving linear maps $\mathbb{R}^{n}\to\mathbb{R}^{n}$. For instance, the symmetry group $S(K_{n})$ of the unit cross-polytope $K_{n}=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}|x_{i}|\leq 1\}$ is the group of order $2^{n}n!$ consisting of signed permutations of the coordinates. Any centrally symmetric convex body $C$ induces a norm on $\mathbb{R}^{n}$ via $\lVert x\rVert_{C}=\inf\{t>0:x\in tC\}$ for $x\in\mathbb{R}^{n}$.
By a kissing configuration of a convex body $C$, we mean a set $S\subset\mathbb{R}^{n}$ such that the translates $x+\frac{1}{2}C,\ x\in S$ have disjoint interiors and each $x+\frac{1}{2}C$ intersects $\frac{1}{2}C$ only at its boundary. Note that $S$ is a kissing configuration of $C$ if and only if it is a kissing configuration of the central symmetrization $\frac{1}{2}(C-C)\in\mathcal{K}_{n}^{0}$, so without loss of generality we assume henceforth that $C\in\mathcal{K}_{n}^{0}$. Under this assumption, a kissing configuration is equivalently a set $S\subseteq\mathbb{R}^{n}$ such that $\lVert x\rVert_{C}=1$ for all $x\in S$ and $\lVert x-x^{\prime}\rVert_{C}\geq 1$ for all $x,x^{\prime}\in S$, $x\neq x^{\prime}$.
The maximum size of a kissing configuration is denoted $\kappa_{T}(C)$ and sometimes called the Hadwiger number of $C$. The maximum size of a kissing configuration of $C$ consisting of the shortest non-zero vectors (with respect to the norm $\lVert\cdot\rVert_{C}$) of a lattice is denoted $\kappa_{L}(C)$. Recall that a lattice $\Lambda$ is the $\mathbb{Z}$-span of a basis of $\mathbb{R}^{n}$. Clearly, $\kappa_{L}(C)\leq\kappa_{T}(C)$. The condition that the shortest non-zero vectors of a lattice $\Lambda$ form a kissing configuration is equivalent to the condition that $\min_{x\in\Lambda\setminus\{0\}}\lVert x\rVert_{C}=1$, and the condition that the kissing configuration is as large as possible is equivalent to the set $\mathcal{M}_{C}(\Lambda)=\{x\in\Lambda:\lVert x\rVert_{C}=1\}$ having maximal cardinality.
We call two lattices $\Lambda_{1}$ and $\Lambda_{2}$ $C$-equivalent, if $\Lambda_{1}=\sigma(\Lambda_{2})$ for some $\sigma\in\mathcal{G}_{C}$. Note that $|\mathcal{M}_{C}(\Lambda)|$ depends only on the $C$-equivalence class of $\Lambda$. Define the support of a vector $x=(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$ to be the set $\text{supp}(x)=\{i:x_{i}\neq 0\}$. We say that $i$ is a max-coordinate of $x$ if $|x_{i}|=\lVert x\rVert_{\infty}$. We define the set $D_{n}^{+}\subset\mathbb{R}^{n}$ to be the union of $D_{n}$ and its translate $D_{n}+(\frac{1}{2},\dots,\frac{1}{2})$, where $D_{n}=\{x\in\mathbb{Z}^{n}:\sum_{i=1}^{n}x_{i}\equiv 0\pmod{2}\}$. It is easily seen that $D_{n}^{+}$ is a lattice if and only if $n$ is even. For instance, $D_{8}^{+}$ is isometric to the $E_{8}$ lattice, while $D_{4}^{+}$ is a scaled Hadamard rotation of the integer lattice $\mathbb{Z}^{4}$. We define the ($\ell^{1}$-norm) covering radius of a discrete set $S\subset\mathbb{R}^{n}$ to be the quantity $\rho(S)=\sup_{y\in\mathbb{R}^{n}}\inf_{x\in S}\lVert x-y\rVert_{1}$. We call the points $y\in\mathbb{R}^{n}$ achieving the supremum deep holes of $S$.
3 The lattice kissing number of $K_{n}$
In this section, we focus on the lattice kissing number problem for $K_{n}$, and prove the following theorem:
Theorem 2.
$\kappa_{L}(K_{4})=40$, achieved precisely by lattices $K_{4}$-equivalent to $D_{4}^{+}$.
Proof.
Let $L\subset\mathbb{R}^{4}$ be a lattice achieving $\kappa_{L}(K_{4})$, and let $\mathcal{M}$ be a set containing exactly one of each $\pm$ pairs of vectors in $\mathcal{M}_{K_{4}}(L)$. Suppose that $|\mathcal{M}|\geq 20$. Note that if $x,y\in\mathcal{M}$ are distinct vectors such that $x\equiv y\pmod{2L}$, then $\frac{x\pm y}{2}\in\mathcal{M}_{K_{4}}(L)$ and the triangle inequality implies that $\text{supp}(x)\cap\text{supp}(y)=\emptyset$, and in particular $|\text{supp}(x)|\leq 2$ or $|\text{supp}(y)|\leq 2$.
Since there are $152^{4}-1=$ 15 non-zero classes in $L/2L$, the following is true: $\mathcal{M}$ contains $k\in\{0,1,2,3,4\}$ vectors whose support has size one and $5-k$ pairs $\{x_{1}^{i},x_{2}^{i}\}_{i=1}^{5-k}$ of vectors such that $x_{1}^{i}\equiv x_{2}^{i}\pmod{2L}$ and $|\text{supp}(x_{1}^{i})|=|\text{supp}(x_{2}^{i})|=2$ for all $i=1,\dots,5-k$. Note that there must exist at least one pair $\{x_{1}^{i_{1}},x_{2}^{i_{1}}\}$ satisfying $\lVert x_{1}^{i_{1}}\rVert_{\infty}=\lVert x_{2}^{i_{1}}\rVert_{\infty}=\frac{1}{2}$. Indeed, if $x,y\in\mathcal{M}$ share a max-coordinate, then $\lVert x\rVert_{\infty}\leq\frac{1}{2}$ or $\lVert y\rVert_{\infty}\leq\frac{1}{2}$, and every $x\in\mathcal{M}$ with $|\text{supp}(x)|>1$ has $|x_{i}|\leq\frac{1}{2}$ for all $i\in\cup_{x\in\mathcal{M}:|\text{supp}(x)|=1}\text{supp}(x)$. If $k\leq 2$, then in addition to the pair $\{x_{1}^{i_{1}},x_{2}^{i_{1}}\}$ there are two vectors $x_{3},x_{4}\in\mathcal{M}$ such that $\lVert x_{3}\rVert_{\infty}=\lVert x_{4}\rVert_{\infty}=\frac{1}{2}$ and $|\text{supp}(x_{3})|=|\text{supp}(x_{4})|=2$. Thus, replacing $L$ by a $K_{4}$-equivalent lattice, we may assume that $L$ contains the lattice $L^{\prime}$ generated by $\mathbb{Z}^{4}$, the vectors $(\frac{1}{2},\frac{1}{2},0,0)$, $(0,0,\frac{1}{2},\frac{1}{2})$ and their midpoint $(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$.
Since $|\mathcal{M}_{K_{4}}(L^{\prime})|=20$ and $\rho(L^{\prime})\leq 1$ by the next lemma, $L$ contains in addition to $L^{\prime}$ at least one of its deep holes. By the next lemma, $L\setminus L^{\prime}$ contains a vector of one of the following shapes.
$v_{1}$ $=(x_{1},\frac{1}{2}-x_{1},x_{2},\frac{1}{2}-x_{2}),\quad 0\leq x_{1},x_{2}\leq\frac{1}{2},$
$v_{2}$ $=(x_{1},x_{1}-\frac{1}{2},x_{2},\frac{1}{2}-x_{2}),\quad 0\leq x_{1},x_{2}\leq\frac{1}{2},$
$v_{2}^{\prime}$ $=(x_{1},\frac{1}{2}-x_{1},x_{2},x_{2}-\frac{1}{2}),\quad 0\leq x_{1},x_{2}\leq\frac{1}{2},$
$v_{3}$ $=(x_{1},x_{1}-\frac{1}{2},x_{2},x_{2}-\frac{1}{2}),\quad 0\leq x_{1},x_{2}\leq\frac{1}{2}.$
Suppose $v_{1}\in L$. Then since
$\lVert(x_{1},\frac{1}{2}-x_{1},x_{2},\frac{1}{2}-x_{2})-(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})\rVert_{1}\leq 1$
for all $0\leq x_{1},x_{2}\leq\frac{1}{2}$ with equality if and only if $x_{1},x_{2}\in\{0,\frac{1}{2}\}$, $L$ contains $(\frac{1}{2},0,\frac{1}{2},0)$ and so $L$ contains $\frac{1}{2}D_{4}^{+}$.
Suppose $v_{2}\in L$. Then
$\lVert(x_{1},x_{1}-\frac{1}{2},x_{2},\frac{1}{2}-x_{2})-(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})\rVert_{1}\leq 1$
for all $\frac{1}{4}\leq x_{1}\leq\frac{1}{2}$, $0\leq x_{2}\leq\frac{1}{2}$ with equality if and only if $x_{2}\in\{0,\frac{1}{2}\}$. Similarly,
$\lVert(x_{1},x_{1}-\frac{1}{2},x_{2},\frac{1}{2}-x_{2})-(\frac{1}{4},-\frac{3}{4},\frac{1}{4},\frac{1}{4})\rVert_{1}\leq 1$
for all $0\leq x_{1}\leq\frac{1}{4}$, $0\leq x_{2}\leq\frac{1}{2}$ with equality if and only if $x_{2}\in\{0,\frac{1}{2}\}$. Thus, $L$ contains $(x_{1},x_{1}-\frac{1}{2},0,\frac{1}{2})$ or $(x_{1},x_{1}-\frac{1}{2},\frac{1}{2},0)$ so it contains $(2x_{1}-\frac{1}{2},2x_{1}-\frac{1}{2},0,0)$. Since $-\frac{1}{2}\leq 2x_{1}-\frac{1}{2}\leq\frac{1}{2}$, we must have $x_{1}\in\{0,\frac{1}{4},\frac{1}{2}\}$. Thus, $L$ contains $(\frac{1}{2},0,\frac{1}{2},0)$ or $(\frac{1}{4},-\frac{1}{4},0,\frac{1}{2})$ so it contains $\frac{1}{2}D_{4}^{+}$ or the lattice $L_{0}$ spanned by $L^{\prime}$ and $(\frac{1}{4},-\frac{1}{4},0,\frac{1}{2})$. The assumption $v_{2}^{\prime}\in L$ leads to the same conclusion.
Now suppose $v_{3}\in L$. Then $2v_{3}+(0,1,0,1)=(2x_{1},2x_{1},2x_{2},2x_{2})\in L$. It is easy to see that this implies $(2x_{1},2x_{1},2x_{2},2x_{2})\in L^{\prime}$. Then the only possibilities are $x_{1},x_{2}\in\{0,\frac{1}{4},\frac{1}{2}\}$ or $x_{1},x_{2}\in\{\frac{1}{8},\frac{3}{8}\}$. In the former case, $L$ contains $\frac{1}{2}D_{4}^{+}$ or $L_{0}$. In the latter case, $L$ contains the lattice $L_{1}$ spanned by $L^{\prime}$ and $(\frac{1}{8},-\frac{3}{8},\frac{1}{8},-\frac{3}{8})$ or the lattice $L_{1}^{\prime}$ spanned by $L^{\prime}$ and $(\frac{1}{8},-\frac{3}{8},\frac{3}{8},-\frac{1}{8})$. Note that $L_{1}$ is $K_{4}$-equivalent to $L_{1}^{\prime}$.
In conclusion, $L$ contains $\frac{1}{2}D_{4}^{+}$, $L_{0}$, $L_{1}$ or $L_{1}^{\prime}$. Suppose that $L$ strictly contains one of these lattices. Then by the reasoning above, $L$ contains at least two of these lattices. This however implies that $\min_{x\in L\setminus\{0\}}\lVert x\rVert_{1}<1$, a contradiction. In conclusion, $L\in\{\frac{1}{2}D_{4}^{+},L_{0},L_{1},L_{1}^{\prime}\}$. We check that
$|\mathcal{M}_{K_{4}}(\frac{1}{2}D_{4}^{+})|=40,\quad|\mathcal{M}_{K_{4}}(L_{0})|=36,\quad|\mathcal{M}_{K_{4}}(L_{1})|=|\mathcal{M}_{K_{4}}(L_{1}^{\prime})|=28,$
and conclude that $L=\frac{1}{2}D_{4}^{+}$.
Lemma 1.
Let $\mathcal{H}_{2}$ be the planar lattice generated by $(1,0)$ and $(\frac{1}{2},\frac{1}{2})$. Then the set of deep holes of the lattice $\mathcal{H}_{2}\oplus\mathcal{H}_{2}\subset\mathbb{R}^{4}$ is the set
$\left\{(x_{1},\pm(\frac{1}{2}-x_{1}),x_{2},\pm(\frac{1}{2}-x_{2})):0\leq x_{1},x_{2}\leq\frac{1}{2}\right\}+\mathcal{H}_{2}\oplus\mathcal{H}_{2}.$
In particular, $\rho(\mathcal{H}_{2}\oplus\mathcal{H}_{2})=1$.
Proof.
The deep holes of a direct sum $L_{1}\oplus L_{2}$ of lattices $L_{1}$, $L_{2}$ are given by $(x,y)$, where $x$ is a deep hole of $L_{1}$ and $y$ is a deep hole of $L_{2}$. Thus, it suffices to show that the deep holes of $\mathcal{H}_{2}$ consist of the translates by vectors in $\mathcal{H}_{2}$ of the line segments $\{(x,\pm(\frac{1}{2}-x)):0\leq x\leq\frac{1}{2}\}$. This is clear. ∎
In a recent article [13, Theorem 3.5], the authors prove that $\kappa_{L}(B_{n}^{p})=O(\frac{n}{p}e^{\frac{n}{p}})$ for all $0<p\leq 2$, where $B_{n}^{p}=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}|x_{i}|^{p}\leq 1\}$ is the unit $\ell^{p}$-ball. The next theorem improves this to $O(2^{n})$ for the $\ell^{1}$-ball, which is similar to the upper bound $\kappa_{L}(C_{s})\leq 2(2^{n}-1)$ which holds for strictly convex $C_{s}\in\mathcal{K}_{n}^{0}$.
Theorem 3.
$\kappa_{L}(K_{n})<12(2^{n}-1)$ for all $n\geq 1$.
Proof.
Let $L\subset\mathbb{R}^{n}$ be a lattice and $\mathcal{M}$ a set of representatives of $\pm$ pairs of vectors in $\mathcal{M}_{K_{n}}(L)$. Suppose that $|\mathcal{M}|\geq 6(2^{n}-1)$. As in the proof of Theorem 2, this implies that $\mathcal{M}$ contains $\ell\geq 2^{n}-1$ vectors whose support has size at most $\lfloor\frac{n}{6}\rfloor$. But the number of vectors whose support has size $k\leq n$ in any kissing configuration in $\mathbb{R}^{n}$ is at most $\binom{n}{k}3^{k}$ by Hadwiger’s bound. Thus, $\ell\leq\sum_{k=1}^{\lfloor\frac{n}{6}\rfloor}\binom{n}{k}3^{k}\leq\lfloor\frac{n}{6}\rfloor\binom{n}{\lfloor\frac{n}{6}\rfloor}3^{\lfloor\frac{n}{6}\rfloor}$, which contradicts $\ell\geq 2^{n}-1$ by routine estimation. ∎
4 Asymptotic bounds for $\kappa_{T}(K_{n})$
4.1 An asymptotic upper bound for $\kappa_{T}(K_{n})$
In this section, we turn our attention from the lattice kissing number problem to the translative kissing number problem, which is known to be a more challenging problem. The idea behind Theorem 1 is that a maximum sized ball inscribed in a cross-polytope need only be enlargened by a factor of approximately $r=\frac{2e}{\sqrt{2\pi e}}\approx 1.315$ (for large $n$) in order for it to have a volume equal to that of the cross-polytope; or more precisely, $\text{vol}(K_{n})\sim\frac{1}{\sqrt{2}}\text{vol}(\frac{r}{\sqrt{n}}B_{n})$ as $n\to\infty$. Recall that $\text{vol}(K_{n})=\frac{2^{n}}{n!}$.
Thus, if the inscribed balls are enlargened by a factor of $\sqrt{2}$, and if most of the balls are contained in a cross-polytope of relatively small radius, then a volume comparison should give a better bound than $13^{n}-$ 1. Here we will use the fact that it is relatively simple to bound the volume of a sphere outside of a cross-polytope by considering certain spherical caps. Since the enlargened balls need not be disjoint, we attach a density to each ball such that the density at any given point of space does not exceed one, as in Blichfeldt’s method.
In our estimations, the following limit will appear often: for a fixed real number $0\leq a\leq 1$, $\lim_{n\to\infty}\frac{1}{n}\log_{2}\binom{n}{\lfloor an\rfloor}=H\left(a\right)$, where $H(s)=-s\log_{2}(s)-(1-s)\log_{2}(1-s)$ and $H(0)=0=H(1)$ is the binary entropy function defined for $0\leq s\leq 1$.
Proof of Theorem 1.
Let $X=\{x_{1},\dots,x_{N}\}$ be a cross-polytope kissing configuration, so that the interiors of $x_{i}+\frac{1}{2}K_{n}$ and $x_{j}+\frac{1}{2}K_{n}$ are disjoint for all $i\neq j$. Let $X^{\prime}=\{x\in X:\text{$x$ has at least $\lfloor cn\rfloor$ coordinates of absolute value}\geq\frac{b}{n}\}$. We select parameter values $c=0.296$ and $b=0.334$. Note that for $x\in X\setminus X^{\prime}$, the sum of the absolute values of $n-\lfloor cn\rfloor$ smallest coordinates (in absolute value) is at most $b(1-c)+\frac{b}{n}=a$. Thus, if $x,x^{\prime}\in X\setminus X^{\prime}$ are such that the positions of $\lfloor cn\rfloor$ largest coordinates coincide (say positions $i=1,\dots,\lfloor cn\rfloor$), then $\sum_{i=1}^{\lfloor cn\rfloor}|x_{i}-x_{i}^{\prime}|\geq 1-2a$. Hence, the interiors of the $\lfloor cn\rfloor$-dimensional bodies $(x_{1},\dots,x_{\lfloor cn\rfloor})+\frac{1-2a}{2}K_{\lfloor cn\rfloor}$ and $(x_{1}^{\prime},\dots,x_{\lfloor cn\rfloor}^{\prime})+\frac{1-2a}{2}K_{\lfloor cn\rfloor}$ are disjoint, and both are contained in $\frac{3-2a}{2}K_{\lfloor cn\rfloor}$. By a volume comparison, we obtain
$|X\setminus X^{\prime}|$ $\leq\binom{n}{\lfloor cn\rfloor}\left(\frac{3-2a}{1-2a}\right)^{\lfloor cn\rfloor}\leq 2^{H(c)(1+o(1))n}\left(\frac{3-2a}{1-2a}\right)^{\lfloor cn\rfloor}$
$\leq 2.9161^{(1+o(1))n}.$
Next we will bound $M=|X^{\prime}|$ using Blichfeldt’s enlargement method. Note that $\frac{1}{\sqrt{n}}B_{n}\subset K_{n}$, so the interiors of $x_{i}+\frac{1}{2\sqrt{n}}B_{n}$ and $x_{j}+\frac{1}{2\sqrt{n}}B_{n}$ are disjoint for all $i\neq j$. Let $\delta_{n}(r)=1-2nr^{2}$ if $0\leq r\leq\frac{1}{\sqrt{2n}}$ and $\delta_{n}(r)=0$ else. The inequality of Blichfeldt [10] (see also the proof of [12, Theorem 29.2]) yields that for every $y\in\mathbb{R}^{n}$ and every subset $\mathcal{X}\subseteq X^{\prime}$,
$2|\mathcal{X}|\sum_{x\in\mathcal{X}}\lVert y-x\rVert_{2}^{2}\geq\sum_{x,x^{\prime}\in\mathcal{X}}\lVert x-x^{\prime}\rVert_{2}^{2}\geq\frac{|\mathcal{X}|(|\mathcal{X}|-1)}{n},$
which implies that
$\sum_{x\in X^{\prime}}\delta_{n}(\lVert y-x\rVert_{2})=\sum_{x\in\mathcal{X}_{y}}\delta_{n}(\lVert y-x\rVert_{2})$ $\leq 1\qquad\text{for all $y\in\mathbb{R}^{n}$,}$
where $\mathcal{X}_{y}=\{x\in X^{\prime}:\lVert y-x\rVert_{2}\leq\frac{1}{\sqrt{2n}}\}$. Let $R=1.5675$ and consider the following bound for the volume of $RK_{n}$:
$\text{vol}(RK_{n})$ $\geq\int_{RK_{n}}\sum_{x\in X^{\prime}}\delta_{n}(\lVert y-x\rVert_{2})dy$
[…]
$=\frac{2M}{2+n}\text{vol}(\frac{1}{\sqrt{2n}}B_{n})-\sum_{x\in X^{\prime}}\text{vol}((x+\frac{1}{\sqrt{2n}}B_{n})\setminus RK_{n}).$
Let $\{v_{i}\}_{i=1}^{2^{n}}$ be an enumeration of the vectors (±$(\pm 1,\dots,\pm 1)$, and define the half-spaces $\mathcal{H}_{i}=\{y\in\mathbb{R}^{n}:y\cdot v_{i}\geq R\}$, which bound the body $RK_{n}$. Let $x\in X^{\prime}$ be fixed, and note that
$\text{vol}((x+\frac{1}{\sqrt{2n}}B_{n})\setminus RK_{n})\leq\sum_{i=1}^{2^{n}}\text{vol}(\mathcal{H}_{i}\cap(x+\frac{1}{\sqrt{2}n}B_{n})).$
Note that $\mathcal{H}_{i}\cap(x+\frac{1}{\sqrt{2}n}B_{n})$, if non-empty, is a spherical cap of radius $r=\frac{1}{\sqrt{2n}}$ and height $h=r-r_{0}$, where $r_{0}=\frac{R-x\cdot v_{i}}{\sqrt{n}}$ is the distance from $x$ to the hyperplane bounding $\mathcal{H}_{i}$. Recall that $x$ has at least $\lfloor cn\rfloor$ coordinates, say coordinates $1,\dots,\lfloor cn\rfloor$, of absolute value $\geq\frac{b}{n}$. Thus, $x\cdot v_{i}\leq 1-\frac{2kb}{n}$, where $k$ is the number of $j=1,\dots,\lfloor cn\rfloor$ such that $x_{j}\cdot v_{ij}<0$. Consequently, $r_{0}\geq\frac{R-1+\frac{2kb}{n}}{\sqrt{n}}$. Approximating a spherical cap by an appropriate cylinder it is easy to see that $V_{\text{cap}}(r,h)\leq h(r^{2}-(r-h)^{2})^{\frac{n-1}{2}}\text{vol}(B_{n-1})=V_{\text{cyl}}(r,h)$. Recall that $\frac{\text{vol}(B_{n})}{\text{vol}(B_{n-1})}\sim\sqrt{\frac{2\pi}{n}}$. We obtain the upper bound
$\sum_{i=1}^{2^{n}}\text{vol}(\mathcal{H}_{i}\cap(x+\frac{1}{\sqrt{2}n}B_{n}))$ $\leq\sum_{k=0}^{\lfloor\frac{(1+\frac{1}{\sqrt{2}}-R)n}{2b}\rfloor}\binom{\lfloor cn\rfloor}{k}2^{n-\lfloor cn\rfloor}V_{\text{cyl}}(r,r-\frac{R-1+\frac{2kb}{n}}{\sqrt{n}})$
$\leq\alpha^{(1+o(1))n}\text{vol}(\frac{1}{\sqrt{2n}}B_{n}),$
where $\alpha=\sup_{0\leq k^{\prime}\leq\frac{1+\frac{1}{\sqrt{2}}-R}{2b}}2^{cH(\frac{k^{\prime}}{c})+(1-c)}(1-2(R-1+2k^{\prime}b)^{2})^{\frac{1}{2}}<1$.
In conclusion,
$R^{n}\frac{2^{n}}{n!}=$ $\text{vol}(RK_{n})\geq M\left(\frac{2}{2+n}\text{vol}(\frac{1}{\sqrt{2n}}B_{n})-\alpha^{(1+o(1))n}\text{vol}(\frac{1}{\sqrt{2n}}B_{n})\right)$
from which it follows that $M\leq 2.91616^{(1+o(1))n}$. ∎
4.2 An asymptotic lower bound for $\kappa_{T}(K_{n})$
In the other direction, Larman and Zong showed that $\kappa_{T}(K_{n})\geq 1.1249^{(1-o(1))n}$ [8], which was consequently improved to $\kappa_{T}(K_{n})\geq\kappa_{T}(K_{n})\geq 1.1348^{(1-o(1))n}$ by Talata [14]. The latter result is also contained in [15], where the author also provides weaker constructive bounds based on algebraic geometry codes. The main result of this section is stated below.
Theorem 4.
$\kappa_{T}(K_{n})\geq 1.1637^{(1-o(1))n}$ as $n\to\infty$.
Let $m>0$ and let $\mathcal{S}$ be a finite subset of $\partial(mK_{n})$ such that for all $x\in\mathcal{S}$, the cardinality of the set $\mathcal{B}(x)=\{y\in\mathcal{S}:\lVert x-y\rVert_{1}<m\}$ is independent of the center $x$. This is achieved when $\mathcal{S}$ is the set of signed permutations of one vector $v\in\mathbb{R}^{n}$. Then by the union bound,
$\kappa_{T}(K_{n})\geq\frac{|\mathcal{S}|}{|\mathcal{B}(x)|}.$ (1)
Thus, the relevant problem is to choose the set $\mathcal{S}$ such that the probability that two uniformly random $x,x^{\prime}\in\mathcal{S}$ satisfy $\lVert x-x^{\prime}\rVert_{1}<m$, is as small as possible. The results mentioned in the previous paragraph are obtained using the set $\mathcal{S}=\{x\in\{0,\pm 1\}^{n}:\sum_{i=1}^{n}|x_{i}|=m\}$ for a suitable integer $m$. However, since we are working with the $\ell^{1}$-norm, it is possible to decrease the aforementioned probability by replacing the set $\{0,\pm 1\}$ with a larger set. Here we will consider for simplicity the set $\{0,\pm 1,\pm 2\}$, which already gives a noticeable improvement.
Proof.
Let $n>0$ and $m_{1}\geq m_{2}>0$ be integers to be determined later and define the set
$\mathcal{X}_{m_{1},m_{2}}$ =$=\{x\in\{0,\pm 1,\pm 2\}^{n}:|x_{i}|=j\text{ for $m_{j}$ values of $i$ ($j=1,2$)}\}$
of cardinality $|\mathcal{X}_{m_{1},m_{2}}|=\binom{n}{m_{1}}\binom{n-m_{1}}{m_{2}}2^{m_{1}+m_{2}}$. Note that $\lVert x\rVert_{1}=m_{1}+2m_{2}:=m$ for all $x\in\mathcal{X}_{m_{1},m_{2}}$. For $x\in\mathcal{X}_{m_{1},m_{2}}$, let $\mathcal{B}_{m_{1},m_{2}}(x)=\{y\in\mathcal{X}_{m_{1},m_{2}}:\lVert x-y\rVert_{1}<m\}$. Clearly $|\mathcal{B}_{m_{1},m_{2}}(x)|$ does not depend on the center $x$ so we may consider the ball centered at $x_{0}=(1,\dots,1,2,\dots,2,0,\dots,0)\in\mathcal{X}_{m_{1},m_{2}}$. For $x\in\mathcal{X}_{m_{1},m_{2}}$, let
$c_{1}(x)$ =|$=|\{1\leq i\leq m_{1}:x_{i}=2\}|,$
$c_{2}(x)$ =|$=|\{m_{1}+1\leq i\leq m_{1}+m_{2}:x_{i}=2\}|,$
$c_{3}(x)$ =|$=|\{1\leq i\leq m_{1}+m_{2}:x_{i}=1\}|.$
Then $x\in\mathcal{B}_{m_{1},m_{2}}(x_{0})$ if and only if $2c_{1}(x)+4c_{2}(x)+2c_{3}(x)>m$. We obtain the upper bound
$|\mathcal{B}_{m_{1},m_{2}}(x_{0})|$ $\leq\sum_{0\leq x_{1}+x_{2}\leq m_{2}}\binom{m_{1}}{x_{1}}\binom{m_{2}}{x_{2}}\binom{m_{1}+m_{2}-x_{1}-x_{2}}{h(x_{1},x_{2})}$
$\binom{n-x_{1}-x_{2}-h(x_{1},x_{2})}{m_{1}-h(x_{1},x_{2})}\binom{n-m_{1}-x_{1}-x_{2}}{m_{2}-x_{1}-x_{2}}2^{m_{1}+m_{2}-x_{1}-x_{2}-h(x_{1},x_{2})}$
$:=\sum_{0\leq x_{1}+x_{2}\leq m_{2}}g(x_{1},x_{2};m_{1},m_{2}),$
where $h(x_{1},x_{2})=\lfloor\frac{m}{2}-x_{1}-2x_{2}\rfloor+1$ is the smallest integer satisfying $2x_{1}+4x_{2}+2h(x_{1},x_{2})>m$. We will consider $m_{1},m_{2}$ satisfying $m_{2}<\frac{m_{1}}{2}$, whence $0\leq h(x_{1},x_{2})\leq m_{1}$ holds for all admissible $x_{1},x_{2}$.
Now let $m_{i}=\lfloor nz_{i}\rfloor$ and $x_{i}=\lfloor ny_{i}\rfloor$ for $i=1,2$, where $0<z_{1},z_{2}\leq 1$ are fixed parameters satisfying $z_{2}<\frac{z_{1}}{2}$ and $z_{1}+z_{2}<1$, and $y_{1},y_{2}\geq 0$ are variables satisfying $0\leq y_{1}+y_{2}\leq z_{2}$. Define
$f(y_{1},y_{2};z_{1},z_{2})=\lim_{n\to\infty}\frac{1}{n}\log_{2}(g(\lfloor ny_{1}\rfloor,\lfloor ny_{2}\rfloor;\lfloor nz_{1}\rfloor,\lfloor nz_{2}\rfloor)).$
We can now write down an explicit expression for $f$.
$f(y_{1},y_{2};z_{1},z_{2})$
$=z_{1}H\left(\frac{y_{1}}{z_{1}}\right)+z_{2}H\left(\frac{y_{2}}{z_{2}}\right)+(z_{1}+z_{2}-y_{1}-y_{2})H\left(\frac{\frac{z_{1}}{2}+z_{2}-y_{1}-2y_{2}}{z_{1}+z_{2}-y_{1}-y_{2}}\right)$
$+(1+y_{2}-\frac{z_{1}}{2}-z_{2})H\left(\frac{y_{1}+2y_{2}+\frac{z_{1}}{2}-z_{2}}{1+y_{2}-\frac{z_{1}}{2}-z_{2}}\right)$
$+(1-z_{1}-y_{1}-y_{2})H\left(\frac{z_{2}-y_{1}-y_{2}}{1-z_{1}-y_{1}-y_{2}}\right)+\frac{z_{1}}{2}+y_{2}.$
We take $z_{1}^{*}=\frac{19}{100}$ and $z_{2}^{*}=\frac{9}{100}$, and find that $\sup_{0\leq y_{1}+y_{2}\leq z_{2}^{*}}f(y_{1},y_{2};z_{1}^{*},z_{2}^{*})\approx 1.17029$, attained at $(y_{1},y_{2})\approx(0.01728,0.04327)$. We conclude that for $m_{i}^{*}=\lfloor nz_{i}^{*}\rfloor$,
$\liminf_{n\to\infty}$ $\frac{1}{n}\log_{2}\left(\kappa_{T}(K_{n})\right)\geq\lim_{n\to\infty}\frac{1}{n}\log_{2}|\mathcal{X}_{m_{1}^{*},m_{2}^{*}}|-\limsup_{n\to\infty}\frac{1}{n}\log_{2}|\mathcal{B}_{m_{1}^{*},m_{2}^{*}}(x_{0})|$
$\geq H(z_{1}^{*})+(1-z_{1}^{*})H\left(\frac{z_{2}^{*}}{1-z_{1}^{*}}\right)+z_{1}^{*}+z_{2}^{*}-\sup_{0\leq y_{1},y_{2}\leq z_{2}^{*}}f(y_{1},y_{2};z_{1}^{*},z_{2}^{*})$
$\approx 0.218818,$
which shows that $\kappa_{T}(K_{n})\geq 2^{0.21881(1-o(1))n}\geq 1.1637^{(1-o(1))n}$. ∎
It is interesting to note that Theorem 4 gives a better existence result for the kissing number of the cross-polytope than the best known existence result $\kappa_{T}(B_{n})\geq 2^{0.2075(1+o(1))n}$ for the $\ell^{2}$-ball mentioned in the introduction [5].